10000 ‘Roshani’ light bulbs with a mean life of 150 days are installed in a new factory and their length of life is normally distributed with standard deviation of 10 days. How many Roshani-bulbs will expire in less than 160 days? 


1
Expert's answer
2021-06-07T17:17:24-0400

"P(X<160)=P(Z<\frac{160-150}{10})=P(Z<1)=0.8413."

Number of Roshani-bulbs which will expire in less than 160 days:

"n=10000*0.8413=8413."